Three people are working on one project. The first person is in charge of 1/4 of the project and the second person is in charge of 2/5. How much of the project the third person is in charge of? Thanks.
Answer by jorel1380(3719)

So, what is left after 1/4 and 2/5? Then:
1 -1/4 -2/5=20/20- 5/20- 8/20=7/20ths of the job left for the 3rd person. ☺☺☺☺